A job order cost sheet for Ryan Company is shown below.

Job No. 92

For 2,000 Units


Date

Direct
Materials

Direct
Labor

Manufacturing
Overhead

Beg. bal. Jan. 1 6,400 6,560 4,592
8 6,560
12 10,000 7,500
25 2,800
27 4,000 3,000
15,760 20,560 15,092
Cost of completed job:
   Direct materials $15,760
   Direct labor 20,560
   Manufacturing overhead 15,092
Total cost $51,412
Unit cost ($51,412 ÷ 2,000) $25.71


(a) On the basis of the foregoing data, answer the following questions.

(1) What was the balance in Work in Process Inventory on January 1 if this was the only unfinished job?

Balance in Work in Process Inventory on January 1 $

Entry field with correct answer17552


(2) If manufacturing overhead is applied on the basis of direct labor cost, what overhead rate was used in each year? (Round answers to 0 decimal places, e.g. 55%.)

Last year Current year
Overhead rate

        Entry field with correct answer70

%

Entry field with incorrect answer

%      What is the current year %?


(b) Prepare summary entries at January 31 to record the current year’s transactions pertaining to Job No. 92. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1
Concepts and reason

Cost accounting is an accounting process in which the costs involved in the process, project are recorded and analyzed. The purpose of cost accounting is to allocate costs to different products and thereby determine the selling price per unit.

Work in Progress: Work in progress in cost accounting implies a value of goods whose production has not yet been completed. Work in progress is the stage of production between raw material and the finished goods. Work in progress includes the cost of material, labor and overhead expenses applied at a different stage of production.

Fundamentals

Direct materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Direct labor: Workers or employees directly involved in the manufacture of goods and services are called direct labor. Direct labor can be directly traceable to a specific product or cost center.

Predetermined overhead rate: It is that rate based on which the overhead expenses are absorbed over the period. It is generally a blanket rate which is further multiplied with the most important factor of production which can be labor hours or machine hours to calculate the overhead expense for the year.

(a) (1)

Calculate opening WIP using the equation as follows:

OpeningWIP=Directmaterial+Directlabor+Manufacturingoverhead=$6,400+$6,560+4,592=$17,552\begin{array}{c}\\{\rm{Opening WIP}} = {\rm{Direct material}} + {\rm{Direct labor}} + {\rm{Manufacturing overhead}}\\\\ = \$ 6,400 + \$ 6,560 + 4,592\\\\ = \$ 17,552\\\end{array}

Hence, opening WIP on January 1 is $17,552.

[Part a1]

Part a1

Part a1

(a) (2)

Compute the manufacturing overhead rate using MS Excel as follows:

Computation of Overhead Rate
Current Year
Particulars
Manufacturing Overhead
Direct Labor
Overhead Rate
Last Year
$10,500
$4,

Hence, the overhead rate as a percentage of labor cost in the previous year is 70% and 75% in the current year.

Working Notes:

Compute the total manufacturing overheads in the current year using the equation as follows:

ManufacturingOverheads=(Overheadsincurredon12Jan+Overheadsincurredon27Jan)=$7,500+$3,000=$10,500\begin{array}{c}\\{\rm{Manufacturing Overheads}} = \left( \begin{array}{l}\\{\rm{Overheads incurred on 12 Jan}} + \\\\{\rm{Overheads incurred on 27 Jan}}\\\end{array} \right)\\\\ = \$ 7,500 + \$ 3,000\\\\ = \$ 10,500\\\end{array}

Hence, the manufacturing overheads in the current year are $10,500.

Compute the total direct labor in the current year using the equation as follows:

DirectLabor=(Laborexpensesincurredon12Jan+Laborexpensesincurredon27Jan)=$10,000+$4,000=$14,000\begin{array}{c}\\{\rm{Direct Labor}} = \left( \begin{array}{l}\\{\rm{Labor expenses incurred on 12 Jan}} + \\\\{\rm{Labor expenses incurred on 27 Jan}}\\\end{array} \right)\\\\ = \$ 10,000 + \$ 4,000\\\\ = \$ 14,000\\\end{array}

Hence, the direct labor expenses incurred in the current year are $14,000.

Compute the overhead rate for the previous year using the equation as follows:

Overheadrate=ManufacturingoverheadDirectlabor×100=$4,592$6,560×100=70%\begin{array}{c}\\{\rm{Overhead rate}} = \frac{{{\rm{Manufacturing overhead}}}}{{{\rm{Direct labor}}}} \times 100\\\\ = \frac{{\$ 4,592}}{{\$ 6,560}} \times 100\\\\ = 70\% \\\end{array}

Hence, the overhead rate in the previous year is 70%.

Compute the overhead rate for the current year using the equation as follows:

Overheadrate=ManufacturingoverheadDirectlabor×100=$10,500$14,000×100=75%\begin{array}{c}\\{\rm{Overhead rate}} = \frac{{{\rm{Manufacturing overhead}}}}{{{\rm{Direct labor}}}} \times 100\\\\ = \frac{{\$ 10,500}}{{\$ 14,000}} \times 100\\\\ = 75\% \\\end{array}

Hence, the overhead rate in the current year is 75%.

[Part a2]

Part a2

Part a2

(b)

Pass a journal entry to record transfer of direct material expense to WIP inventory using MS Excel as follows:

Post
Date
Description
Debit(S)
Credit(S)
ref
Work in Progress Inventory
Direct Materials Inventory
S9.360
31-Jan
$9,360
|(To

Working Note:

Compute the direct materials inventory using the equation as follows:

Directmaterials=(Directmaterialpurchasedon8Jan+Directmaterialpurchasedon25Jan)=$6,560+$2,800=$9,360\begin{array}{c}\\{\rm{Direct materials}} = \left( \begin{array}{l}\\{\rm{Direct material purchased on 8 Jan}}\\\\ + {\rm{Direct material purchased on 25 Jan}}\\\end{array} \right)\\\\ = \$ 6,560 + \$ 2,800\\\\ = \$ 9,360\\\end{array}

Hence, the direct materials purchased in January are $9,360.

[Part b]

Part b

Part b

Pass a journal entry to record transfer of wages payable to WIP inventory using MS Excel as follows:

Post
Debit(S)
Date
Description
Credit(S)
ref.
Work in Progress Inventory
Direct labor wages payable
(To record transfer of wa

Working Note:

Compute the direct labor wages payable using the equation as follows:

Directlaborwages=(Directlaborexpenseson12Jan+Directmaterialpurchasedon27Jan)=$10,000+$4,000=$14,000\begin{array}{c}\\{\rm{Direct labor wages}} = \left( \begin{array}{l}\\{\rm{Direct labor expenses on 12 Jan}}\\\\ + {\rm{Direct material purchased on 27 Jan}}\\\end{array} \right)\\\\ = \$ 10,000 + \$ 4,000\\\\ = \$ 14,000\\\end{array}

Hence, the direct materials purchased in January are $14,000.

Pass a journal entry to record transfer of manufacturing overheads to WIP Inventory using MS Excel as follows:

Post
Description
Debit(S)
Credit(S)
Date
ref.
Work in Progress Inventory
Manufacturing Overheads
(To record transfer of manuf

Working Note:

Compute the total manufacturing overheads using the equation as follows:

ManufacturingOverheads=(Manufacturingoverheadsexpenseson12Jan+Manufacturingoverheadsexpenseson27Jan)=$7,500+$3,000=$10,500\begin{array}{c}\\{\rm{Manufacturing Overheads}} = \left( \begin{array}{l}\\{\rm{Manufacturing overheads expenses on 12 Jan}}\\\\ + {\rm{Manufacturing overheads expenses on 27 Jan}}\\\end{array} \right)\\\\ = \$ 7,500 + \$ 3,000\\\\ = \$ 10,500\\\end{array}

Hence, the manufacturing overheads expenditure in January is $10,500.

Pass a journal entry to record the transfer of finished goods from WIP inventory to finished goods inventory using MS Excel as follows:

Post
Description
Credit(S)
Date
Debit(S)
ref
Finished goods inventory
Work in progress inventory
|(To record transfer of fini

Working Note:

Compute the total work in progress inventory using the equation as follows:

Workinprogressinventory=(Totaldirectmaterial+Totaldirectlabor+Totalmanufacturingoverheads)=$15,760+$20,560+15,092=$51,412\begin{array}{c}\\{\rm{Work in progress inventory}} = \left( \begin{array}{l}\\{\rm{Total direct material}} + {\rm{Total direct labor}}\\\\ + {\rm{Total manufacturing overheads}}\\\end{array} \right)\\\\ = \$ 15,760 + \$ 20,560 + 15,092\\\\ = \$ 51,412\\\end{array}

Ans: Part a1

Opening WIP on January 1 is $17,552

Part a2

The overhead rate as a percentage of labor cost in the previous year is 70% and 75% in the current year.

Part b

Post
Date
Description
Debit(S)
Credit(S)
ref
Work in Progress Inventory
Direct Materials Inventory
S9.360
31-Jan
$9,360
|(To

Post
Debit(S)
Date
Description
Credit(S)
ref.
Work in Progress Inventory
Direct labor wages payable
(To record transfer of wa

Post
Description
Debit(S)
Credit(S)
Date
ref.
Work in Progress Inventory
Manufacturing Overheads
(To record transfer of manuf

Post
Description
Credit(S)
Date
Debit(S)
ref
Finished goods inventory
Work in progress inventory
|(To record transfer of fini